OpenStudy (anonymous):

What must be added to 3x -7 to make x 2 + 4x -1?

OpenStudy (jango_in_dtown):

let y be added to 3x-7 to make x^2+4x-1 then y+(3x-7)=x^2+4x-1 this imply y=x^2+4x-1-3x+7=x^2+x+6 hence the answer is x^2+x+6

OpenStudy (jango_in_dtown):

@kbriones0

OpenStudy (anonymous):

Thank you so much!!